OpenStudy (anonymous):

A remote-control car races straight down the street at 15 feet per second. Four seconds later, a second remote-control car races straight down the same street at 25 feet per second. How long will it take the second car to catch up to the first car? on what scale is the speed of each car measured

OpenStudy (anonymous):

Are there any answers to choose from?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, could you type them in please? :)

OpenStudy (anonymous):

sure hold on

OpenStudy (anonymous):

Ok. :)

OpenStudy (anonymous):

a.) tens b.) hundreds c.) thousands d.) ones

OpenStudy (anonymous):

15 x 4 = 60 feet (total head start) 60 / 25 = 2.4 sec (total head start / 2nd car speed) 15 x 2.4 = 36 feet (additional feet gained by 1st car while 2nd car makes up 60 feet) 36 + 60 = 96 feet (additonal feet + headstart) 96 / 25 = 3.84 sec (Total feet / 2nd car speed) 3.84 + 4 = 7.84 (additional time + headstart) So, with that, what do you think it is? :)

OpenStudy (anonymous):

tens

OpenStudy (anonymous):

I believe that is correct! I mean, it couldn't be ones, there is no such thing. Couldn't be hundredths because it isn't that much so it defiantly couldn't be thousandths. :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Of course! :D